OpenStudy (anonymous):

a triangle has two equal sides and a third side that is five inches more than the length of the other two sides. If the perimeter of the triangle is 32 inches, what is the length of the longest side?

OpenStudy (bahrom7893):

y+y+(5+y) = 32 5+y - ?

OpenStudy (bahrom7893):

5+3y = 32 3y = 27 y = 9

OpenStudy (bahrom7893):

So the longest side is y+5 = 9+5 = 14

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

P = 2x + ( x + 5) = 32 in -> 3x = 32 - 5 = 27 -> x = 27/3 = 9 in => x + 5 = 9 + 5 = 14 in